Lindstrom Company produces two fountain pen models. Information about its products follows: Product A Product B Sales revenue $ 140,000 $ 190,000 Less: Variable costs 56,000 62,700 Contribution margin $ 84,000 $ 127,300 Total units sold 5,000 5,000 Lindstrom’s fixed costs total $85,000. Required: 1. Determine Lindstrom’s weighted-average unit contribution margin and weighted-average contribution margin ratio. (Round your weighted-average CM to 2 decimal places and your CM ratio to 1 decimal place (i.e. .123 should be entered as 12.3%)). 2. Calculate Lindstrom’s break-even point in units and in sales revenue. (Round your "Sales Revenue" answer to 2 decimal places and "Sales Units" answer to the nearest whole number.) 3. Calculate the number of units that Lindstrom must sell to earn a $120,000 profit. (Round your answer to the nearest whole number.) 4. Calculate Lindstrom’s margin of safety and margin of safety as a percentage of sales if it sells 7,000 total pens. (Round your margin of safety in units to the nearest whole number and your percentage of sales answer to 2 decimal places (i.e. .1234 should be entered as 12.34%))
Answer #1

Product A

Product B

Sales

140,000

190,000

Variable costs

56,000

62,700

Contribution Margin

84,000

127,300

Units Sold

5,000

5,000

Contribution Margin per Unit

16.8

25.46

Weighted average CM = 16.8*5000/10,000 + 25.46*5000/10,000

= $21.13

Weighted average selling price = 33

Weighted average CM Ratio = 21.13/33 = 64.03% i.e. 64.0%

2.Break even point in unit sales = Fixed costs/Weighted average contribution margin per unit

= 85,000/21.13

= 4,022.72 units i.e. 4,023 units

Break even sales revenue = Fixed costs/Weighted average CM Ratio

= 85,000/64%

= $132,812.5

3.Target Profit = $120,000

Add: Fixed costs = 85,000

Desired contribution margin = $205,000

Number of units required to be sold = 205,000/21.13

= 9,701.85 units or 9,702 units

4.Margin of safety in units = total sales – break even sales

= 7,000-4,023

= 2,977 units

% = 2,977/7,000 = 42.53%